OpenStudy (anonymous):

Last one... What is the solution to the equation? x + 12 - 20 = 14 + 8 x=? I got 14

OpenStudy (agl202):

Simplifying x + 12 + -20 = 14 + 8 Reorder the terms: 12 + -20 + x = 14 + 8 Combine like terms: 12 + -20 = -8 -8 + x = 14 + 8 Combine like terms: 14 + 8 = 22 -8 + x = 22 Solving -8 + x = 22 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '8' to each side of the equation. -8 + 8 + x = 22 + 8 Combine like terms: -8 + 8 = 0 0 + x = 22 + 8 x = 22 + 8 Combine like terms: 22 + 8 = 30 x = 30 Simplifying x = 30

OpenStudy (anonymous):

oh lol I was way off I'm horrible at math Thank You, you helped out a lot

OpenStudy (agl202):

:D

OpenStudy (anonymous):

I got 100% thank you :D

OpenStudy (agl202):

Yaya!!! That great! Proud 4 ya! :)

OpenStudy (anonymous):

Lol